Lark corporation is considering the acquisition of an asset for use in its business over the next five years. However, Lark must decide whether it would be better served by leasing the asset or buying it. An appropriate asset could be purchased for $15,000.00 and it would qualify as a three-year asset under MACRS classification. Assume that the election to expense assets under section 179 is made, but any available additional first-year depreciation is not claimed, and that the asset is not expected to to have a salvage value at the end of its use by Lark. Alternatively, Lark could lease the asset for a $3,625.00 annual cost over the 5-year period. If Lark is in the 21% tax bracket, would you recommend that Lark buy or lease the asset? In your calculations, assume that 10% is an appropriate discount factor. Please show calculations.
Answer #1

There will be two options to choose from: Purchase the asset or lease the asset.

Option 1: Purchase the asset

Outflow at Point 0 (i.e. Today on the date of purchase) = $ 15,000

Tax Savings for the first year after purchase (i.e. At the end of current year) = $ 15,000 * 100% * 21% = $ 3,150 [Since, the deduction is claimed under section 179 for the whole depreciation in the current year]

Present value of Tax savings = $ 3,150 / [(1 + 0.10)1] = $ 2,863.63

Net Cash Outflow in Purchase Option = $ 15,000 - $ 2,863.63

Net Cash outfow (Option 1) = $ 12,136.36

Option 2: Lease the asset

Annual cost of leasing (Outflow at the year end) = $3,625

Marginal Tax Saving on the lease expense (Tax savings at the year end) = $3,625 * 21% = $ 761.25

Net Cash Outflow = $ 3,625- $761.25 = $ 2,863.75

Present Value of all the Cash outflow in Leasing option = Annual Cash Outflow * Cumulative Annuity factors for five year

= $ 2,863.75 * Cdf at 10% for 5 years

= $ 2,863.75 * 3.79

PV fof Cash Outflow (Option 2) = $ 10,853.61

Final Recommendation:

Since the cash outflow is low in Option 2, we should recomment leasing the asset for the mentioned period of fuve years.